14119 € was deposited on 4.4.2006 in a bank account. The interest rate was 3,5 % (p.a.) and the tax at source rate was 30 %. The bank added the accrued interest to the capital at the end of each year. Calculate the accumulated value of the investment on 4.4.2010 when the account was closed.


1
Expert's answer
2022-02-09T15:46:18-0500

Solution:

Reasoning: Due to calculation of accumulated interest, we are therefore required to find compound interest.

Time from 4.4.2006 - 4.4.2010 = 4 years

For compound Interest, A=P(1+r÷100)n

=14119(1+3.5÷100)4

=14119×1.1475

=16201.5525

I=Amount- principal

=16201.5525 - 14119

I =2082.5525

Tax at the source =30÷100×2082.5525

=625.76575

Accrued Interest = 2082.5525-625.76575

=1457.78675

Accumulated Amount = 14119+1457.78675

= 15576.78675